OpenStudy (anonymous):

The reaction A---> products is first order in A. Initially, [A] = .800M and after 54 min [A] = .100M. At what time is[A] .025M What is the rate of reaction when [A] is .025M I got 267.35 min and 1.94 x 10^-6 Ms^-1 can someone please tell me if this is correct

OpenStudy (aaronq):

ln[A]=-kt + ln[Ao] ln(0.1) = -k(54 min) + ln(0.8) k= 0.0385/min ln(0.025) = 0.0385t + ln(0.8) t= 90.2 min i haven't done these in awhile but i think thats right

OpenStudy (aaronq):

oh wait you wanted rate lol sorry r=k[A] r=0.0358/min(0.025)=0.000895/min